# Large-Scale Calculations of $\beta$-Decay Rates and Implications for $r$-Process Nucleosynthesis

### v1

@@ Abstract
Nuclear $\beta$ decay is a key element of the astrophysical rapid neutron capture process ($r$-process). In this paper, we present state-of-the-art global $\beta$-decay calculations based on the quantified relativistic nuclear energy density functional theory and the deformed proton-neutron quasiparticle random-phase approximation. Our analysis considers contributions from allowed and first-forbidden transitions. We used two point-coupling functionals with carefully calibrated time-odd terms and isoscalar pairing strength. The new calculations display consistent results for both employed functionals, especially near the neutron drip line, suggesting slower $\beta$ decays past the $N=126$ neutron shell closure than in commonly used $\beta$-decay models. The new rates, along with the existing rates based on the latest non-relativistic calculations, are found to slow down the synthesis of heavy elements in the $r$-process and significantly reduce the contribution of neutron-induced fission.
1 Introduction
@@
The rapid neutron capture process ($r$-process) is responsible for synthesizing the heaviest elements that naturally exist in the Universe (Burbidge et al., 1957; Cameron, 1957; Horowitz et al., 2019; Cowan et al., 2021; Arcones & Thielemann, 2022). Although its exact site remains elusive, it is expected to occur in explosive and neutron-rich astrophysical events, such as compact binary mergers and rare types of core-collapse supernovae. The process involves thousands of species of neutron-rich nuclei, many of which have yet to be experimentally studied. They undergo various nuclear reactions and decays, including neutron capture, photodissociation, $\beta$ decay, and fission. As most of the nuclei involved are beyond the current reach of experiments, theoretical description of their properties has a profound impact on our understanding of the $r$-process. As a step towards consistent evaluation of the required nuclear physics input with controlled extrapolation, this work focuses on the influence of $\beta$ decay. The basic role of $\beta$ decay in the $r$-process is to transfer the material from the isotopic chain with proton number $Z$ to the one with $Z+1$. During the $(n,\gamma)\leftrightarrow(\gamma,n)$ equilibrium, the most abundant isotope in a given isotopic chain, so-called the “waiting-point” nucleus, effectively determines the $\beta$-decay rate of the isotopic chain. The half-lives of waiting-point nuclei with neutron magic numbers $N=82$ and $126$ qualitatively explain the origin of the second and third $r$-process abundance peaks at mass numbers $A\sim 130$ and $196$, respectively. After the $(n,\gamma)\leftrightarrow(\gamma,n)$ equilibrium breaks down, $\beta$ decay, as well as $\beta$-delayed neutron emission, neutron capture, and fission, are the main mechanisms that affect the final abundance pattern (Surman et al., 1997; Arcones & Martínez-Pinedo, 2011; Mumpower et al., 2012a, b, c; Goriely et al., 2013; Eichler et al., 2015). Fission determines the end point of the $r$-process path, past $Z\sim 90$ (Panov & Thielemann, 2003; Martínez-Pinedo et al., 2007; Panov, I. V. et al., 2010; Eichler et al., 2015; Giuliani et al., 2018; Mumpower et al., 2018; Vassh et al., 2019; Giuliani et al., 2020), redistributing and increasing the relative abundance of heavy nuclei. This mechanism is referred to as “fission (re)cycling”. Neutron-induced fission has been discussed as the dominant fission channel (Panov et al., 2005; Martínez-Pinedo et al., 2007; Petermann et al., 2012); however, depending on the time scale and amount of the population of fissile nuclei, the contribution of neutron-induced fission may be significantly altered.
@@
In the majority of $r$-process studies, $\beta$-decay data have been taken from the phenomenological Finite Range Droplet Model (FRDM), combined with the quasiparticle random phase approximation (QRPA) (Möller et al., 1997, 2003, 2019). More microscopic approaches, whether based on the nuclear shell model (Sharma et al., 2022; De Gregorio et al., 2024) or ab initio methods (Gysbers et al., 2019), remain limited due to the unfavorable scaling of problem complexity with system size. Leveraging the balance between scalability and rigorous microscopic description, the energy density functional (EDF) theory is a suitable many-body method for providing reliable predictions of nuclear properties for $r$-process studies, as it is computationally tractable across the entire chart of nuclides. Starting from the nuclear ground state, excitations can be obtained by considering the linear response of the nucleus within QRPA. Until recently, most large-scale calculations of $\beta$-decay rates based on the nuclear EDF framework and QRPA have been restricted to spherical nuclei (Engel et al., 1999; Niu et al., 2013; Marketin et al., 2016). Only a single global calculation has considered axially-deformed nuclei (Ney et al., 2020). Since global sets of theoretical $\beta$-decay rates are scarce, the literature on the sensitivity of $r$-process simulations on $\beta$ decay has been limited (Niu et al., 2013; Mumpower et al., 2014; Lund et al., 2023; Kullmann et al., 2023; Hao et al., 2023), with the primary focus on the $r$-process observables such as the abundance pattern and electromagnetic emission. The dynamic effect of $\beta$ decay on the evolution of other reactions, decays, and abundances has rarely been explicitly discussed. In particular, the impact of $\beta$-decay modeling on the fission contribution has been largely unexplored.
@@
In this Letter, we present new global $\beta$-decay calculations using the recently developed axially-deformed relativistic QRPA framework Ravlić et al. (2024), with the residual interaction optimized in Ravlić & Nazarewicz (2025). Using this tool, we investigate the effect of the $\beta$ decay on the $r$-process dynamics, through comparisons with other global models.
2 Method
2.1 Calculation of $\beta$-decay half-lives
@@
The ground-state of the nucleus is obtained with the relativistic Hartree-Bogoliubov theory assuming axially-deformed and reflection-invariant nuclear shapes (Nikšić et al., 2014). Two point-coupling interactions are employed in this work: DD-PC1 (Nikšić et al., 2008) and DD-PCX (Yüksel et al., 2019), both adopting separable pairing interaction (Tian et al., 2009). The ground-state quadrupole deformation is obtained by performing constrained calculations with the quadrupole deformation $\beta_{2}\in[-0.5,0.5]$, and subsequently releasing the constraint to find the configuration that minimizes the total energy. Odd-$A$ and odd-odd nuclei are treated by self-consistent quasiparticle blocking within the equal-filling approximation (Perez-Martin & Robledo, 2008). Calculations are performed by discretizing quasiparticle states in a stretched axially-deformed harmonic oscillator basis. We use $N_{\rm osc}=18$ shells for nuclei up to $Z=103$ and for heavier nuclei $N_{\rm osc}=20$. To obtain excited states in the daughter nucleus, we employ the linear response approach, developed in Ravlić et al. (2024). The residual interaction couplings, including the isoscalar pairing interaction strength $V_{0}^{is}$, and Landau-Migdal coupling $g_{0}$, in addition to effective axial-vector coupling $g_{A}$, are used as optimized in the recent work (Ravlić & Nazarewicz, 2025). The $\beta$-decay rates $\lambda_{\beta}$ are calculated by the contour integration around a suitably chosen contour $\mathcal{C}$ encircling the $Q_{\beta}$ window:
@@
where $\lambda_{n(p)}$ is the neutron(proton) chemical potential, and $\Delta_{np}=1.293$ MeV is the neutron-proton mass difference. The nuclear matrix elements are contained within the shape-factor
@@
where $k,ka,kb,$ and $kc$ are tabulated in Behrens & Bühring (1971); Bambynek et al. (1977). The super-allowed decay constant is taken as $K=6137$ s. The circular contour is discretized with a 30-point Gauss-Laguerre quadrature. The integration over electron energy $W$ is performed by Lagrange interpolation of the integrand on a 20-point Chebyshev grid, as discussed in Ravlić & Nazarewicz (2025). Subsequently, the half-lives are obtained as $T_{1/2}=\text{ln}2/\lambda_{\beta}$. We include the allowed Gamow-Teller (GT) $1^{+}$ transitions, as well as the first-forbidden (FF) $0^{-},1^{-}$, and $2^{-}$ $\beta$ decays. Calculations are performed in the range of nuclei between $Z=20$ and $Z=110$, containing 4598 individual nuclides. For each isotopic chain, calculations start close to the valley of stability, and end at the two-neutron drip-line theoretically estimated in Ravlić et al. (2023).
Table 1: Summary of global $\beta$-decay models considered in this work for $r$-process simulations, including the method used to compute the rates, restrictions on nuclear shapes, and contribution of different multipoles. GT corresponds to $1^{+}$ and $F$ to $0^{-},1^{-}$ and $2^{-}$ multipoles.
@@
FRDM Möller et al. (2019) micro-macro + QRPA(schematic) deformed GT+FF(gross)
D3C Marketin et al. (2016) EDF(covariant) + QRPA(matrix) spherical GT+$F$
SkO Ney et al. (2020) EDF(Skyrme) + QRPA(linear response) deformed GT+$F$
@@
2.2 Nuclear reaction network
@@
The $r$-process abundance evolution is calculated with the nuclear reaction network code PRISM (Sprouse, 2020). In addition to our new DD-PC1 and DD-PCX sets of $\beta$-decay rates, a summary of $\beta$-decay models employed in this work is shown in Table 1. In particular, it includes the FRDM + QRPA $\beta$-decay tables from Möller et al. (2003, 2019), based on a schematic separable QRPA interaction and phenomenological estimates of FF transitions. Other rates based on the EDF framework include the predictions by Ney et al. (2020), calculated with the axially-deformed charge-exchange linear response QRPA with the Skyrme SkO’ functional, and relativistic matrix-QRPA results with the EDF D3C* from Marketin et al. (2016), restricted to spherical nuclei. Other nuclear properties, such as one-neutron separation energies and fission data, are primarily based on the FRDM model and experimental data. Separate calculations were also performed with the REACLIB nuclear data library (Cyburt et al., 2010). See the Supplementary Material for the details and results (Ravlić et al., 2025).
@@
For astrophysical conditions, we first employ the parameterized wind model of Panov & Janka (2009), with an initial entropy per baryon of $s/k_{B}=40$, and an expansion timescale of 20 ms. The initial electron fraction $Y_{e}$, a measure of the neutron-richness of the environment, ranges from 0.15 to 0.35 in 0.025 intervals. The initial temperature of these astrophysical trajectories is 10 GK, and the initial compositions of the nuclei were calculated using the SFHo equation of state (Steiner et al., 2013). Additionally, we also use a trajectory with an initial condition with $s/k_{B}=20$ and $Y_{e}=0.05$, to simulate a neutron-rich environment in which fission cycling is expected to occur.
@@
To investigate the effect of the new $\beta$-decay rates in more realistic astrophysical conditions, we consider thermodynamical evolution from two sets of astrophysical simulations: a magneto-rotational supernova (Reichert et al., 2021a; Obergaulinger & Aloy, 2017; Reichert et al., 2021b), in which it has been suggested that the production of nuclei past the third $r$-process peak at $A\sim 195$ occurs; and outflow from the accretion disk formed after a merger of binary neutron stars (Metzger & Fernández, 2014; Holmbeck et al., 2019). Similarly to the case of the parametrized trajectories, nucleosynthesis calculations were initiated at $T=10$ GK, and the composition at the time was calculated using the SFHo equation of state.
3 Results and Discussion
@@
Figure 1: Comparison between (a) mean averaged half-lives $\bar{r}$ and (b) their spread $\sigma$ obtained in this work ($D$-PC1 and $D$-PCX) with the results of SkO’ model (Ney et al., 2020), D3C* (Marketin et al., 2016), and FRDM (Möller et al., 2003). The experimental data are taken from  (Kondev et al., 2021). Averaging in Eq. (5) is performed over the nuclei with half-lives up to the maximum experimental half-life $T_{1/2}^{\rm exp,max}$ in the range from $10^{-1}$ s to $10^{4}$ s.
@@
Figure 2: Comparison of theoretical $\beta$-decay rates relative to predictions of our $D$-PC1 EDF model: (a) our $D$-PCX model, (b) SkO’ model (Ney et al., 2020), (c) D3C* model (Marketin et al., 2016), and (d) FRDM model (Möller et al., 2019).
3.1 Benchmarking theoretical rate calculations

### v2

@@ Abstract
Nuclear $\beta$ decay is a key element of the astrophysical rapid neutron capture process ($r$-process). In this work, we present state-of-the-art global $\beta$-decay calculations based on the quantified relativistic nuclear energy density functional theory and the deformed proton-neutron quasiparticle random-phase approximation. Our analysis considers contributions from allowed and first-forbidden transitions. We used two point-coupling functionals with carefully calibrated time-odd terms and isoscalar pairing strength. The new calculations display consistent results for both employed functionals, especially near the neutron drip line, suggesting slower $\beta$ decays past the $N=126$ neutron shell closure than in commonly used $\beta$-decay models. The new rates, along with the existing rates based on the recent non-relativistic global calculations, are found to slow down the synthesis of heavy elements in the $r$-process and significantly reduce the contribution of neutron-induced fission.
I Introduction
@@
The rapid neutron capture process ($r$-process) is responsible for synthesizing the heaviest elements that naturally exist in the Universe [BBFH, Cameron, Horowitz2019, Cowan2021, Arcones2022]. Although its exact site remains elusive, it is expected to occur in explosive and neutron-rich astrophysical events, such as compact binary mergers and rare types of core-collapse supernovae. The process involves thousands of species of neutron-rich nuclei, many of which have yet to be experimentally studied. They undergo various nuclear reactions and decays, including neutron capture, photodissociation, $\beta$ decay, and fission. As most of the nuclei involved are beyond the current reach of experiments, theoretical description of their properties has a profound impact on our understanding of the $r$-process. As a step towards consistent evaluation of the required nuclear physics input with controlled extrapolation, this work focuses on the influence of $\beta$ decay. The basic role of $\beta$ decay in the $r$-process is to transfer the material from the isotopic chain with proton number $Z$ to the one with $Z+1$. During the $(n,\gamma)\leftrightarrow(\gamma,n)$ equilibrium, the most abundant isotope in a given isotopic chain, so-called the “waiting-point” nucleus, effectively determines the $\beta$-decay rate of the isotopic chain. The half-lives of waiting-point nuclei with neutron magic numbers $N=82$ and $126$ qualitatively explain the origin of the second and third $r$-process abundance peaks at mass numbers $A\sim 130$ and $196$, respectively. After the $(n,\gamma)\leftrightarrow(\gamma,n)$ equilibrium breaks down, $\beta$ decay, as well as $\beta$-delayed neutron emission, neutron capture, and fission, are the main mechanisms that affect the final abundance pattern [Surman_rep_1997, Arcones_rep_2011a, Mumpower_rep_2012a, Mumpower_rep_2012b, Mumpower_rep_2012c, Goriely2013, Eichler2015]. Fission determines the end point of the $r$-process path, past $Z\sim 90$ [Panov2003, Martinezpinedo2007, Panov2010, Eichler2015, Giuliani2018, Mumpower2018, Vassh2019, Giuliani2020], redistributing and increasing the relative abundance of heavy nuclei. This mechanism is referred to as “fission (re)cycling”. Neutron-induced fission has been discussed as the dominant fission channel [Panov2005, Martinezpinedo2007, Petermann2012]; however, depending on the time scale and amount of the population of fissile nuclei, the contribution of neutron-induced fission may be significantly altered.
@@
In the majority of $r$-process studies, $\beta$-decay data have been taken from the phenomenological Finite Range Droplet Model (FRDM), combined with the quasiparticle random phase approximation (QRPA) [Moller1997, Moller2003, Moller2019]. More microscopic approaches, whether based on the nuclear shell model [Sharma2022, DeGregorio2024] or ab initio methods [Gysbers2019, Li2025], remain limited due to the unfavorable scaling of problem complexity with system size. Leveraging the balance between scalability and rigorous microscopic description, the energy density functional (EDF) theory is a suitable many-body method for providing reliable predictions of nuclear properties for $r$-process studies, as it is computationally tractable across the entire chart of nuclides. Starting from the nuclear ground state, excitations can be obtained by considering the linear response of the nucleus within QRPA. Extending the model space beyond two-quasiparticle configurations included in QRPA, although leading to improved agreement with experiment, remains numerically challenging and was applied to specific nuclei only [Robin2016, Robin2018a, Liu2024a]. Until recently, most large-scale calculations of $\beta$-decay rates based on the nuclear EDF framework and QRPA have been restricted to spherical nuclei [Engel1999a, Niu2013, Marketin_RQRPA_2016]. Only a single global calculation has considered axially-deformed nuclei [Ney2020]. Since global sets of theoretical $\beta$-decay rates are scarce, the literature on the sensitivity of $r$-process simulations on $\beta$ decay has been limited [Niu2013, Mumpower2014, Lund2023, Kullmann2023, Hao2023], with the primary focus on the $r$-process observables such as the abundance pattern and electromagnetic emission. The dynamic effect of $\beta$ decay on the evolution of other reactions, decays, and abundances has rarely been explicitly discussed. In particular, the impact of $\beta$-decay modeling on the fission contribution has been largely unexplored.
@@
In this work, we present new global $\beta$-decay calculations using the recently developed axially-deformed relativistic QRPA framework [Ravlic2024a], with the residual interaction optimized in Ref. [Ravlic2025a]. Using this tool, we investigate the effect of the $\beta$ decay on the $r$-process dynamics, through comparisons with other global models.
This paper is outlined as follows. In Sec. II we describe theoretical methods for both nuclear $\beta$ decay and $r$-process calculations. Section III presents benchmarks of $\beta$-decay rates with experimental data and other theoretical calculations, discussions on $Q_{\beta}$-values and impact of first-forbidden transitions, as well as quantified impact of the new rates on different $r$-process scenarios. Conclusions and an outlook are given in Sec. IV. All results compiled in this work are available in a public repository [GitHub], together with an online interactive visualization tool to display observables across the nuclide chart [rpx_website].
II Theoretical Methods
II.1 Calculation of $\beta$-decay half-lives
@@
The ground-state of the nucleus was obtained with the relativistic Hartree-Bogoliubov theory assuming axially-deformed and reflection-invariant nuclear shapes [Niksic2014a]. Two point-coupling interactions were employed in this work: DD-PC1 [Niksic2008a] and DD-PCX [Yuksel2019a], both adopting separable pairing interaction [Tian2009]. The ground-state quadrupole deformation was obtained by performing constrained calculations with the quadrupole deformation $\beta_{2}\in[-0.5,0.5]$, and subsequently releasing the constraint to find the configuration that minimizes the total energy. Odd-$A$ and odd-odd nuclei were treated by self-consistent quasiparticle blocking within the equal-filling approximation (EFA) [Perez-Martin2008a]. Calculations were performed by discretizing quasiparticle states in a stretched axially-deformed harmonic oscillator basis. We used $N_{\rm osc}=18$ shells for nuclei up to $Z=103$ and for heavier nuclei $N_{\rm osc}=20$. To obtain excited states in the daughter nucleus, we employed the linear response approach, developed in Ref. [Ravlic2024a]. The residual interaction couplings, including the isoscalar pairing interaction strength $V_{0}^{is}$, and Landau-Migdal coupling $g_{0}$, in addition to effective axial-vector coupling $g_{A}$, were used as optimized in the recent work [Ravlic2025a]. The $\beta$-decay rates $\lambda_{\beta}$ were calculated by the contour integration around a suitably chosen contour $\mathcal{C}$ encircling the $Q_{\beta}$ window:
@@
where $\lambda_{n(p)}$ is the neutron(proton) chemical potential, and $\Delta_{np}=1.293$ MeV is the neutron-proton mass difference. The nuclear matrix elements are contained within the shape factor
@@
where $k,ka,kb,$ and $kc$ are tabulated in Refs. [Behrens1971a, Bambynek1977a]. The super-allowed decay constant was taken as $K=6137$ s. The circular contour was discretized with a 30-point Gauss-Laguerre quadrature. The integration over electron energy $W$ was performed by Lagrange interpolation of the integrand on a 20-point Chebyshev grid, as discussed in Ref. [Ravlic2025a]. Subsequently, the half-lives were obtained as $T_{1/2}=\text{ln}2/\lambda_{\beta}$. We included the allowed Gamow-Teller (GT) $1^{+}$ transitions, as well as the first-forbidden (FF) $0^{-},1^{-}$, and $2^{-}$ $\beta$ decays. Calculations were performed in the range of nuclei between $Z=20$ and $Z=110$, containing 4598 individual nuclides. For each isotopic chain, calculations started close to the valley of stability, and ended at the two-neutron drip line theoretically estimated in Ref. [Ravlic2023a].
Table 1: Summary of global $\beta$-decay models considered in this work for $r$-process simulations, including the method used to compute the rates, restrictions on nuclear shapes, and contribution of different multipoles. GT corresponds to $1^{+}$ and FF to $0^{-},1^{-}$ and $2^{-}$ multipoles.
@@
FRDM [Moller2019] micro-macro + QRPA(schematic) deformed GT+FF(gross)
D3C [Marketin_RQRPA_2016] EDF(covariant) + QRPA(matrix) spherical GT+$F$
SkO [Ney2020] EDF(Skyrme) + QRPA(linear response) deformed GT+$F$
@@
II.2 Nuclear reaction network
@@
The $r$-process abundance evolution was calculated with the nuclear reaction network code PRISM [SprouseTheis2020]. In addition to our new DD-PC1 and DD-PCX sets of $\beta$-decay rates, a summary of $\beta$-decay models employed in this work is shown in Table 1. In particular, it includes the FRDM + QRPA $\beta$-decay tables from Refs. [Moller2003, Moller2019], based on a schematic separable QRPA interaction and phenomenological estimates of FF transitions. Other rates based on the EDF framework include the predictions from Ref. [Ney2020], calculated with the axially-deformed charge-exchange linear response QRPA with the Skyrme SkO’ functional, and relativistic matrix-QRPA results with the EDF D3C* from Ref. [Marketin_RQRPA_2016], restricted to spherical nuclei. Other nuclear properties, such as $\beta$-delayed neutron emission probabilities, one-neutron separation energies, and fission data, are primarily based on the FRDM model and experimental data. Separate calculations were also performed with the REACLIB nuclear data library [Cyburt2010]. See Sec. II.3 for more details.
@@
For astrophysical conditions, we first employed the parameterized wind model of Ref. [Panov2009], with an initial entropy per baryon of $s/k_{B}=40$, and an expansion timescale of 20 ms. The initial electron fraction $Y_{e}$, a measure of the neutron-richness of the environment, ranges from 0.15 to 0.35 in 0.025 intervals. The initial temperature of these astrophysical trajectories is 10 GK, and the initial compositions of the nuclei were calculated using the SFHo equation of state [Steiner2013]. Additionally, we also used a trajectory with an initial condition with $s/k_{B}=20$ and $Y_{e}=0.05$, to simulate a neutron-rich environment in which fission cycling is expected to occur.
@@
To investigate the effect of the new $\beta$-decay rates in more realistic astrophysical conditions, we considered thermodynamical evolution from two sets of astrophysical simulations: a magneto-rotational supernova [Reichert2021, Obergaulinger2017, Reichert_github], in which it has been suggested that the production of nuclei past the third $r$-process peak at $A\sim 195$ occurs; and outflow from the accretion disk formed after a merger of binary neutron stars [Metzger2014, Holmbeck2019]. Similarly to the case of the parametrized trajectories, nucleosynthesis calculations were initiated at $T=10$ GK, and the composition at the time was calculated using the SFHo equation of state.
II.3 Nuclear Data
We employed two sets of nuclear data libraries for reactions and decays other than $\beta$ decay: the FRDM-based nuclear data library, which is described in the following and also in, e.g., Refs. [Holmbeck2019, saito2025]; and the REACLIB nuclear data library [Cyburt2010]. To simplify the calculation, for both sets of nuclear data, we adopted symmetric fission yields without neutron emission.
In the FRDM-based nuclear data library, neutron capture rates were calculated using the statistical Hauser-Feshbach (HF) code CoH3 with nuclear masses from FRDM2012 [Moller2016] or AME2020 [Wang2021] whenever experimental data are available. Photodissociation rates, the reverse reaction of neutron capture, were calculated within PRISM using the detailed balance based on the neutron capture rates and the one-neutron separation energies. Theoretical $\alpha$-decay rates were estimated using the Viola–Seaborg relation [Viola1966] including the $Q_{\alpha}$ values based on FRDM2012 and AME2020. Whenever possible, these were replaced with experimental $\alpha$-decay rates from NUBASE2020 [Nubase2020]. Spontaneous fission rates were taken from NUBASE2020. Neutron-induced fission rates were calculated with CoH3 using the masses from FRDM2012 or AME2020 and the fission barriers from FRLDM [Moller2015]. $\beta$-delayed fission rates as well as neutron-emission probabilities were calculated using the FRDM+QRPA+HF framework [Mumpower2016, Mumpower2018].
For the calculations with the REACLIB nuclear data library, most reaction and decay rates, including the photodissociation rates, were taken directly from REACLIB. However, spontaneous fission and $\beta$-delayed fission rates were taken from NUBASE2020. For the latter, the rates were supplemented by FRDM-based rates [Mumpower2018]. For neutron-induced fission, we used data from Ref. [Panov2010] calculated with FRDM masses [Moller1995] to maintain consistency with other REACLIB data. For both nuclear data sets, theoretical rates were replaced by the data from NUBASE2020 whenever available.
III Results and Discussion
@@
Figure 1: Global comparison between (a) mean averaged half-lives $\bar{r}$ and (b) their spread $\sigma$ obtained in this work ($D$-PC1 and $D$-PCX) with the results of SkO’ model [Ney2020], D3C* [Marketin_RQRPA_2016], and FRDM [Moller2003]. The experimental data are taken from Ref. [Nubase2020]. Averaging in Eq. (5) is performed over the nuclei with half-lives up to the maximum experimental half-life $T_{1/2}^{\rm exp,max}$ in the range from $10^{-1}$ s to $10^{4}$ s.
@@
Figure 2: Comparison of theoretical $\beta$-decay rates relative to predictions of our $D$-PC1 EDF model: (a) our $D$-PCX model, (b) SkO’ model [Ney2020], (c) D3C* model [Marketin_RQRPA_2016], and (d) FRDM model [Moller2019].
III.1 Benchmarking theoretical rate calculations